Write down the equation of the graph of the exponential function of the form y=a^x which passes through the point (1,2).
The co-ordinates specified are (1,2) and (1,5)
I thought (1,6)=6=ab^1
(2,9) 9=ab^2
9=6b^2
root of 9/6=root 6/2
However; it seems the correct answer is listed as y=6^x Where am I going wrong?

Answer by rothauserc(4718)

y = a^x, then use logarithm rule
x = log(base a)y, and a = y
these curves intersect at (0,1)
indeed for the point (1,6)
y = 6^x
likewise
y = 2^x for (1,2)
y = 5^x for (1,5)